### v1

# Perception of Visual Content: Differences Between Humans and Foundation Models

Abstract
Human-annotated content is often used to train machine learning (ML) models. However, recently, language and multi-modal foundational models have been used to replace and scale-up human annotator’s efforts. This study compares human-generated and ML-generated annotations of images representing diverse socio-economic contexts. We aim to understand differences in perception and identify potential biases in content interpretation. Our dataset comprises images of people from various geographical regions and income levels washing their hands. We compare human and ML-generated annotations semantically and evaluate their impact on predictive models. Our results show low similarity between human and machine annotations from a low-level perspective, i.e., types of words that appear and sentence structures, but are alike in how similar or dissimilar they perceive images across different regions. Additionally, human annotations resulted in best overall and most balanced region classification performance on the class level, while ML Objects and ML Captions performed best for income regression. Humans and machines’ similarity in their lack of bias when perceiving images highlights how they are more alike than what was initially perceived. The superior and fairer performance of using human annotations for region classification and machine annotations for income regression show how important the quality of the images and the discriminative features in the annotations are.
Introduction
Machine Learning (ML) is prevalent in numerous applications, e.g., facial recognition and recommendation systems. However, many of these use cases could potentially result in discrimination and unfair decisions. For example, it is possible that a facial recognition algorithm is less likely to perform well when predicting people of certain ethnic groups, e.g., people of African heritage may not be identified as accurately as people of European origin. This is problematic, especially as we aim to rely more on ML for daily tasks. Since supervised ML models require data for training, biased training data could affect the fairness and bias of the learned model – also known as algorithmic bias – which can cause it to generate prejudiced predictions. Unfortunately, it is quite likely that the data will contain bias considering that they are commonly annotated by human annotators (Demartini, Roitero, and Mizzaro 2024).
Furthermore, generative models have started to be used to produce labelled data for training and evaluation of ML models, which amplifies this issue (Thomas et al. 2023). It is also unclear how these machine annotations are different to human annotations. Since human annotation could contain biases, one possible approach to creating potentially less biased annotations is to use machine-generated annotations. These annotations could likely be more objective compared to those generated by humans.
In this study, we first investigate how similar annotations generated by humans and machines are, and explore possible justifications for their similarities or differences, in the context of how they think and perceive images. This leads to our first research question (RQ1): “How similar are annotations generated by humans compared to the annotations generated by ML models?” Moreover, due to the aforementioned biases that could appear in the annotations, we look into how different annotations affect fairness in predictive models. This leads to our second research question (RQ2): “How do different combinations of annotations affect fairness in ML predictive models?” To do this, we use the annotations in vector representation to train predictive ML models. These annotations are based on images of people washing their hands from varying geographical regions and socio-economic classes.
This study contributes to the fields of information retrieval and algorithmic bias in several ways. Firstly, it provides a comparative analysis of human-generated and machine-generated annotations in the context of socio-economic diverse content, which give insights on how humans and machines are alike in their bias when perceiving visual content despite their low-level differences, filling a gap in the existing literature. Secondly, it explores whether different annotations sets can impact the efficacy and fairness of ensemble-based ML models, thus advancing both theoretical and practical understanding of this critical issue.
Related Work
Effect of Bias on ML Models. It is inevitable that human annotations for training ML models may contain bias and stereotypes. Numerous qualitative and quantitative studies have demonstrated that there exist biases in numerous techniques, datasets, and models related to Natural Language Processing and Information Retrieval (Baeza-Yates 2020; Bolukbasi et al. 2016; Caliskan, Bryson, and Narayanan 2017). These biases could reinforce stereotypes that already exist in society (Rekabsaz, Kopeinik, and Schedl 2021). Fan et al. (2022) investigated this by conducting a study in which participants were asked to annotate video content associated with various regions and levels of income. When sentiment analysis was performed on the annotations, it was found that they contained regional and income-level biases. Unfortunately, as discussed in their paper, one of the main concerns of these biased annotations is that they impact ML models when used as training data. This causes various search and recommendation systems to use these algorithms to feed users with stereotypical content. Additionally, crowd annotators may be experts in different domains, so high-quality annotations are not always guaranteed. Not only could noise be introduced, but the annotated data may not be consistent (Cao et al. 2023). These unreliable annotations may also introduce more biases, which would affect the ML model’s performance. Devillers, Vidrascu, and Lamel (2005) emphasise how these biases can be dangerous when the training dataset involves actual humans as these learned biases tend to mimic forms of discrimination common amongst humans, such as racial and gender bias. An example of this is Microsoft’s “Tay”, an AI chatbot that initially mimicked a teenage girl but became anti-semitic, racist, and sexist over a span of only 16 hours (Zemčík 2020). Sun, Nasraoui, and Shafto (2020) further discuss how the problem worsens when users make decisions based on the biased content provided by these models. These decisions could then result in data used to train other algorithms in the future, thereby creating a never-ending bias-reinforcement cycle. In order to mitigate this issue and break the loop, more diverse data is needed (Fan et al. 2022). Chhikara et al. (2023); Dash et al. (2019) further support this by discussing how textual content written by numerous different authors will contain a greater variation of perspectives, especially on polarising topics, such as religion and politics. This makes it more crucial to see how fairly different perspectives are reflected when attempting to summarise these into one text. van Atteveldt, van der Velden, and Boukes (2021) looked at comparing sentiment analysis methods in news headlines. They used various approaches to compare the sentiment scores produced for these headlines, including manual annotation, crowd-coding, dictionary approaches, and ML algorithms. In this case, their ML algorithm was used to evaluate the news headlines rather than create equivalent annotations to compare against the human annotations. As compared to this body of literature, we focus on comparing human and machine-generated annotations of visual content with the goal of understanding how bias may differ between these annotations.
Dealing with Imbalanced Data. It is important for data to be well-balanced, ensuring that every class in the dataset is well-represented. This is especially true for traditional ML models as they were originally designed for datasets that are balanced. There are various ML methods that were designed to deal with imbalanced datasets, such as random under-sampling and over-sampling techniques (Estabrooks, Jo, and Japkowicz 2004; Chawla et al. 2002), creating synthetic data (Chen et al. 2023; He et al. 2008; Han, Wang, and Mao 2005), and cost-sensitive learning where the model is penalised more when a minority class sample is incorrectly classified compared to a majority class sample (Ting 2002). In our work, we adopt under-sampling techniques to overcome the issue of data imbalance across some classes in our dataset due to the limited number of available samples.
Ensemble Models. Minimising the trade-off between accuracy and fairness is a question that is very relevant to our work. There has been research indicating that a loss in accuracy is inevitable when a model is trained with limitations relating to fairness. Based on Kenfack et al. (2021), we know that compared to individual classifiers, ensemble models have an improved predictive performance, but compromise on fairness. In our work, we include the use of ensemble models in our experiment to test whether there is the same effect on our dataset.
Cognitive Differences in Humans and Machines. To this day, human experts are still needed to provide accurate and reliable annotations that could be used to train ML models (Valizadegan, Nguyen, and Hauskrecht 2013). Understanding how differently humans and machines think is crucial in helping us develop better and fairer automated systems that could assist us in solving problems, i.e., annotating images for ML. Sharifi Noorian et al. (2022) proposed Scalpel-HS, a framework that is intended to identify unknown unknowns in images by combining human intelligence and machine learning. Unknown unknowns refer to instances that an ML model fails to identify or has not been trained on. This framework aims to reduce cognitive load for humans by using ML models to first identify relevant objects and relationships, so that the humans can focus on evaluating relevance rather than performing the task from scratch. In our study, instead of using ML to aid humans in the annotating process, we investigate how similar human annotations and machine-generated annotations are and factors that explain their similarity or dissimilarity considering how humans and machines perceive images.
How machines think inevitably affects their decision-making. In the context of image recognition tasks, Zhang et al. (2021) investigates how humans and ML models select features differently. Their study revealed that humans are not always better than ML models at selecting discriminative features for image recognition and that humans rely on contextual information, which may affect how they select the features. In our work, instead of image segmentation tasks, we look at how humans and machines perform annotation tasks on images. From this, we explore how well the discriminative features in the annotation sets are by using them to train predictive models.
Methodology
This section details the methodology we used to address the research questions introduced earlier. The methodology begins with data collection from Dollar Street. Annotations are preprocessed and added to the data either through crowdsourcing or ML techniques. The methodology is divided into two key experiments: the first evaluates annotation similarity; the second investigates the impact of different annotation sets on supervised models, using ensemble learning and comparing the performance of human and machine annotations.
Experimental Set-Up
In this study, we utilised CPU resources to conduct all experiments. The computations were performed locally on a device with an Apple M3 Pro Chip, 11 CPU cores, and 36GB of RAM.
Dataset
The dataset utilised in our study was obtained from Dollar Street, a project by Gapminder, and is available under a Creative Commons license CC BY 4.0 (Gapminder 2021). Dollar Street visualises global public data by arranging families based on their income levels, from the poorest on the left to the wealthiest on the right. It features photographs of everyday objects from more than 240 families in 50 countries, providing a representation of living conditions across different economic classes. There are 197 unique instances of images and videos utilised in our experiments, consisting of people washing their hands, sourced from various geographical regions and income levels. The topic of hand washing was selected due to its crucial role in public health and its practice across all cultures and income levels.
Data Annotation
Figure 1: Handwashing images annotated using the pre-trained Faster R-CNN model.
Crowdsourcing. The first annotation set is sourced from a structured crowdsourcing process conducted on Amazon Mechanical Turk (MTurk) by Fan et al. (2022). In their study, crowd workers viewed short videos of individuals washing their hands from various global regions and income levels, providing textual annotations. These annotations captured various aspects of handwashing behaviour and the surrounding environmental context. The dataset includes annotations for 112 videos, each annotated by 10 US-based workers. Fan et al. (2022) implemented quality control measures to ensure reliable annotations. Consent from the participants was obtained through a signed consent form. Additionally, no personally identifiable information or offensive content were collected in the crowdsourcing process. We use these existing human annotations in our current study, benefiting from their comprehensive approach to data collection.
ML Object Detection. Images used in this study are retrieved frame-by-frame from videos capturing handwashing practices. The images are processed using a pre-trained Faster R-CNN model with a ResNet-50-FPN backbone, as outlined by Li et al. (2021), which offers improvements over the original model configuration introduced in Ren et al. (2017). This model is available under the MIT License. This model was originally trained on the COCO dataset (Lin et al. 2014) and is utilised to generate our machine-based object annotations, as illustrated in Figure 1. We selected the Faster R-CNN model for its prominence in object detection, its efficiency, and its effectiveness (i.e., mean average precision) on the COCO labels  (Mohammed, Tannouche, and Ounejjar 2022; Ren et al. 2017). We chose to perform object detection on the dataset as most of the images contain objects, i.e., bottles, sinks, instead of subjects performing actions, i.e., person washing hands. Due to this, the COCO dataset was chosen as the most suitable for training as it contains annotations of everyday objects.
ML Captioning. The images are also passed through a pre-trained BLIP model (Li et al. 2022), with Vision Transformer (ViT) base backbone (Dosovitskiy et al. 2020), for image captioning purposes. The model is available under the BSD 3-Clause ”New” or ”Revised” License. The model was originally trained on the COCO captions dataset (Chen et al. 2015), which helps ensure that this and the object detection model are trained on data that are based on the same images. The BLIP model configuration was chosen as it has achieved state-of-the-art results, outperforming other Vision-Language Pre-training (VLP) methods. We chose to perform image captioning as it would likely resemble the annotations generated by humans.
Data Preparation and Preprocessing
The data cleaning process constituted the initial step in preparing our dataset for analysis. In this phase, corrupted images were identified and removed based on criteria including integrity checks and visual inspections. Images that did not contain labels from any of the annotation categories are also removed, i.e., images that did not have any objects detected in it. This process reduced the original dataset to 197 unique data samples with distinct IDs.
The annotations were preprocessed to deal with inconsistencies in the human annotations. These steps involve noise/punctuation removal, conversion to lower case, auto-correcting misspellings, and segmenting strings that are concatenated.
We utilised a sentence embedding model to convert the three types of annotations into sentence embeddings. Instead of using a pre-trained sentence transformer model to generate the embeddings, we trained a sentence transformer model from scratch on all three sets of annotations. We did this as we want the model and representations to be very focused, and due to the limited number of samples, fine-tuning an existing sentence transformer, such as sBERT, resulted in worse performance.
To create the sentence transformer, we used a pre-trained Transformer model, “distilroberta-base”, which is available at the Hugging Face Hub (Sanh et al. 2019), to obtain the word embedding model. We then combined this word embedding model and a mean pooling function over the token embeddings so that it could generate a fixed-size vector as output. Using the sentence transformer model, we generated embeddings with a size of 384 dimensions.
Experiment 1: Investigating Annotation Similarity (RQ1)
To investigate the similarity between human, machine-generated object labels, and machine-generated captions, we utilised the embeddings that have been generated for the three types of annotations across all images.
In the first part of this experiment, we utilised T-Stochastic Neighbor Embedding (T-SNE) to visualise the vector embeddings in two-dimensional space (van der Maaten and Hinton 2008). Doing this helps us analyse the distances of the vectors, and thus the annotations, for every image in the dataset and see how they are distributed.
We then used the sentence transformer model’s built-in similarity function to compute the cosine similarity scores between each possible pair of annotations. Using these scores, we look at the average similarity scores for each pair and see how the pairwise scores are distributed using scatterplots and boxplots.
These experiments are appropriate for addressing RQ1 as t-SNE and computation of similarity scores across pairs allows us to evaluate how similar the annotations are, whereas the pairwise scatter plot helps us derive trends and patterns between annotations, further evaluating their likeness to one another.
Experiment 2: Exploring How Annotations Impact Fairness in Predictive Models (RQ2)
To investigate how the three annotation sets affect predictive models, we use the vector embeddings of every annotation as training data for two supervised learning models.
Before training the models, the dataset was split into training and test sets with a split ratio of 80:20, respectively . The training set was first used to run Grid Search Cross-Validation (CV) with each of the two models to obtain the optimal parameters for each model whilst performing 5-fold cross-validation. Grid Search CV was performed in every independent run.
The first model is an RUS Boost Classifier used to classify each sample on region (Seiffert et al. 2008). This classifier involves the incorporation of the Adaptive Boost algorithm, which is a sequential ensemble technique that can lower a model’s bias and variance, with random under-sampling, which helps deal with data imbalance (Mohammed and Kora 2023; Seiffert et al. 2008). This makes the algorithm ideal for our dataset due to its small size and there not being an equal number of samples from every region. The base estimator used in the model is the Decision Tree Classifier (Swain and Hauska 1977). After obtaining the optimal parameters, i.e., maximum number of estimators where boosting stops and the learning rate, for the model using grid search CV, we trained the classifier using these parameter values. The trained classifier is run 10 times with different random seeds to obtain a mean F1 score for each class (region) and a mean overall weighted F1 score, with their respective standard deviations. This is done to evaluate the model effectiveness for each class and the overall model performance. The model is run multiple times to allow us to evaluate its stability. The Welch’s t-test was performed to compare whether the difference in performance of the best and worst overall-performing models are statistically significant. This test was performed due to the Gaussian distribution of the F1-Scores and varying variances between the models.
The second model uses the Adaboost.R2 algorithm to predict each sample’s income (Drucker 1997). Similar to the region classifier, this algorithm uses a Decision Tree Regressor as its base estimator. Since boosting helps reduce prediction error on the small dataset, this algorithm is suitable for this task. Using grid search CV, we optimised for the number of estimators, learning rate, and the type of loss function used in the Adaboost.R2 algorithm. The model was executed 10 times with different random seeds. The root mean square error (RMSE) was obtained for every run, both for overall and region-based evaluation. The RMSE takes the square root of the mean square error, which performs the measurement using the same units as the target variable, i.e., income level. A regression plot is generated to visualise the distribution of predicted samples by income for the best-performing model. The Kolmogorov-Smirnov test was performed to compare whether the difference in performance of the best and worst overall-performing regression models are statistically significant. This test was performed due to the non-Gaussian distribution of the RMSE scores between the models.
These predictive models are appropriate in addressing RQ2 as it helps us investigate how often each region is correctly identified for a given image, and how well income is predicted for images in a given region. For evaluation purposes, the F1-score is appropriate and sufficient as it is a good metric for evaluating classification performance on imbalanced datasets, as ours is, whereas the RMSE scores are appropriate and sufficient as the measurement being done in the same units as the target variable makes it easier to interpret. Looking at these evaluation metrics for every region helps us determine whether the models perform better for particular regions, giving insight on how fair the model is.
Results and Discussion
RQ1: How similar are human-generated and ML-generated annotations?
Figure 2: t-SNE Visualisation of Sentence Embeddings of Each Label
Annotations Used Mean ± s.d.
ML Objects and ML Captions 0.87 ± 0.05
ML Objects and Human Labels 0.42 ± 0.1
ML Captions and Human Labels 0.45 ± 0.06
Table 1: Mean and standard deviation of similarity scores between each pair of annotations across all samples.
Figure 3: A plot of pairwise relationships between each pair’s similarity score.
Figure 2 illustrates a t-SNE visualisation of the vector embeddings of the three annotation categories for each data sample. The results show how, despite having trained an embedding model from scratch, the embeddings still form into three distinct clusters. Further analysis shows that the human labels cluster is furthest from the ML captions cluster, with the ML object labels cluster in between them.
The clusters of embeddings are due to differences in the annotations. Human annotations are holistic, encapsulating both the objects and the array of actions unfolding in the captured images and videos. On the other hand, the ML-generated object labels are primarily object-centric. The COCO dataset, which the ML object detection model is trained on, does not extend to action recognition, which is not relevant due to the object-oriented images in the dataset, thereby inherently limiting the ML-generated object annotations to object identification. The ML-generated captions, although more descriptive than the object labels, tend to not be as specific as the human annotations since they are more likely to contain easily recognisable actions and objects rather than complex descriptions of people’s emotions or the environment. Additionally, there are apparent syntactic differences between the annotations - the ML object labels consists of lists of objects per image; the ML captions consist of a phrase per image; while the human labels are a list of words and phrases. It is also important to note that a limitation that could have affected this distinct categorisation is that stop word removal and lemmatisation were not performed during preprocessing as to preserve nuances and context that contain relevant information. This causes the sentence structures in the annotations to remain unchanged, resulting in more pronounced categorisations. Considering these differences, the distinct clusters in the visualisation are expected.
However, since t-SNE is only a two-dimensional representation of the data, it would be useful to also look at the similarity scores, generated by the BLIP model’s cosine similarity function, between each pair of annotations across the dataset. This is shown in Table 1. These values give a more accurate and quantitative comparison of the annotation similarities. It can be seen that the ML object labels and the ML captions are the most similar pair of annotations (0.87), followed by ML captions and human labels (0.45), and ML object labels and human labels (0.42). It is important to note that when comparing human labels with ML-generated ones, the similarity scores are relatively low, regardless of the type of ML annotation.
The high similarity of both ML annotation types is a reasonable outcome. Although the annotations were generated by two different pre-trained ML models which were trained with different types of training data, i.e., labels versus image-text pairs, the training data both originated from the COCO dataset and are based on the same set of images. This means that there would be more overlap of objects between the two annotation categories. In contrast, the low similarity values between either ML annotations with the human labels is expected since there may be higher variety of words and phrases that were used by the human annotators. Moreover, when we compare the t-SNE visualisation and the similarity scores, it may seem inconsistent at first glance due to the ML objects and human labels pair having the lowest similarity score despite being seemingly furthest apart visually. However, since t-SNE is a dimensionality reduction technique, there may be features that were lost in the process, resulting in less accurate representations.
To further explore the relationships of these similarity scores, a plot of pairwise relationships is presented in figure 3. The plot illustrates how similarity scores between two annotation categories correlate to another pair’s scores for a given image. The plot is colour-coded by region to allow further analysis of similarity scores by geographical location. The analysis reveals that there are no particular trends or patterns with respect to region. Moreover, when looking at the subplots for ML objects-ML captions pairs, i.e., subplots on first column/row, the similarity scores tend to be relatively high, on average, with the exception of several outliers. This aligns with the previous observation that ML objects and ML captions have the highest average similarity score (0.87). This pair also has the smallest variance out of all pairs, which is also evident when comparing the vertical spread of scores across the subplots in all rows, i.e., the scores of the ML objects-ML captions pair are not spread as much vertically when disregarding the outliers. In contrast, the ML objects-human labels pair, i.e., second row/column, appears to have the highest variance, which also aligns with the standard deviation of the mean similarity score observed in the previous experiment (0.1). It is also shown that the ML objects-ML captions pair has the same outliers as the ML objects-human labels pair (first row, second column), but not with the ML captions-human labels pair (first row, third column). Interestingly, the ML captions-human labels pair (third column/row) seems to have no outliers.
The lack of discernible patterns in the way the scores are distributed across regions suggests that the images have little to no effect on how similar two sets of annotations will be. This could mean that the discriminative features in the images, or lack thereof, are detected equally by both humans and ML models. Therefore, none of the annotation sets describe images from a particular region any better or worse than the other. This also suggests that humans and machines may be similar in terms of how biased they are when perceiving images.
The low variance of similarity scores for the ML objects-ML captions pair suggests that for the majority of images, both annotations consistently contain similar objects, which - similar to the previous discussion on their high similarity scores - is expected due to the origin of the both models’ training data. Similarly, the high variance of scores for the ML objects-human labels pair aligns with expectations. This is due to, not only syntactic differences, i.e., objects vs longer phrases or sentences, but also differences in the way machines and humans interpret visual content, i.e., machines are more likely to capture easily recognisable objects and actions but may fail to capture more complex and nuanced observations, such as environment aesthetics or partially obstructed objects.
Moreover, the fact that the ML objects-ML captions pair has the same outliers as the ML objects-human labels pair only (first row, second column) indicates that for those particular images, the ML object detection model may be detecting objects that do not appear in the corresponding ML captions or human labels, or failing to identify objects that do. This reasoning is consistent with the findings of Zhang et al. (2021), which suggest that deeper neural networks, such as ResNet - the architecture of our object detection model - may struggle more with difficult images. Perhaps there is obstruction or lack of objects in the image, leading to no or incorrect objects being detected in these outliers, resulting in the low similarity scores when compared to their respective ML Captions and human labels. The ML captioning model as well as human annotators have the advantage in that they are able to capture non-object elements, such as actions, e.g., “wash”, or environmental descriptions, e.g., “dirty”.
RQ2: How do different combinations of annotations affect fairness in ML predictive models?
F1-Score (Mean ± Standard Deviation)
Annotations Used Africa Asia Europe The Americas Overall
ML Objects 0.43 ± 0.05 0.45 ± 0.06 0.30 ± 0.09 0.51 ± 0.10 0.42 ± 0.03
ML Captions 0.60 ± 0.05 0.32 ± 0.05 0.57 ± 0.02 0.38 ± 0.06 0.45 ± 0.02
Human Labels 0.62 ± 0.05 0.54 ± 0.13 0.57 ± 0.05 0.51 ± 0.11 0.56 ± 0.07
ML Objects & ML Captions 0.64 ± 0.07 0.50 ± 0.05 0.49 ± 0.06 0.46 ± 0.09 0.53 ± 0.04
ML Objects & Human Labels 0.57 ± 0.05 0.41 ± 0.05 0.57 ± 0.06 0.55 ± 0.10 0.51 ± 0.04
ML Captions & Human Labels 0.60 ± 0.09 0.38 ± 0.05 0.62 ± 0.07 0.51 ± 0.12 0.51 ± 0.03
ML Objects & ML Captions & Human Labels 0.65 ± 0.06 0.46 ± 0.04 0.51 ± 0.06 0.53 ± 0.11 0.53 ± 0.05
Table 2: Comparison of the region classification model performance across different annotation types. The mean and standard deviation of the F1-Score were recorded over 10 independent runs, both for the overall model performance and for each class individually.
RMSE (Mean ± Standard Deviation)
Annotations Used Africa Asia Europe The Americas Overall
ML Obj 256.17 ± 88.2 1161.35 ± 116.23 2467.03 ± 250.75 1142.27 ± 83.42 1341.65 ± 85.56
ML Captions 179.69 ± 54.86 1456.47 ± 132.90 2164.40 ± 99.97 1815.89 ± 59.91 1564.56 ± 48.54
Human 228.35 ± 55.14 1657.22 ± 107.99 2675.59 ± 104.75 1433.51 ± 74.91 1633.25 ± 49.57
ML Obj & ML Captions 168.79 ± 39.57 1223.70 ± 137.91 2111.71 ± 138.83 1191.59 ± 136.96 1279.39 ± 57.66
ML Objects & Human 117.97 ± 11.04 1497.18 ± 86.39 2621.28 ± 219.83 1108.55 ± 85.18 1480.72 ± 78.63
ML Captions & Human 149.31 ± 16.17 1523.02 ± 108.62 2525.53 ± 186.66 1202.37 ± 72.59 1486.67 ± 86.93
Obj & Captions & Human 125.85 ± 8.27 1407.31 ± 54.97 2549.58 ± 135.91 1191.46 ± 82.55 1449.52 ± 45.33
Table 3: Comparison of the income regression model performance across different annotation types. The mean and standard deviation of the RMSE were recorded over 10 independent runs, both for the overall model performance and for each class individually.
This subsection aims to compare the efficacy of various annotation combinations - ML Objects, ML Captions, and Human Labels - to enhance the performance of models devised for geographical region classification and income regression. The performance of the models was appraised using the F1 score, which balances precision and recall, and Root Mean Square Error (RMSE), which measures performance using the same units as the target variable.
Table 2 tabulates the F1-Scores of the region classification ensemble model across various combinations of the three annotation sets: ML Objects, ML Captions, and Human Labels. The mean and standard deviation of the F1-Scores across 10 independent runs were recorded, both for the overall model and for each class (region).
It is discerned that the Human Labels resulted in the highest overall model performance (0.56). The models utilising combinations of the human and ML annotations exhibited poorer overall performance, followed by the ML Captions Model (0.45) and the ML Objects Model (0.42). Compared to other models, the Human Labels model has the smallest range of F1-Scores across all classes, ranging from 0.51 to 0.62. Moreover, when looking at per-class performance, it is revealed that each class best performs with different models. Africa (class) exhibits the highest performance with the model that uses all three annotation sets (0.65), Asia with the Human Labels Model (0.54), Europe with the ML Captions and Human Labels Model (0.62), and the Americas with the ML Objects and Human Labels Model (0.55). One notable observation is that regardless of overall or per-class performance, human labels are always used in the models - none of the ML-only models resulted in superior overall or per-class performance.
The superior overall performance of the Human Labels model is expected. As discussed by Zhang et al. (2021), humans are more likely to use contextual reasoning in their decision-making process. Although the paper talks about this in the context of humans segmenting images, this may similarly extend to annotation tasks, positively impacting the model’s performance in classifying regions. Perhaps it is the case that classifying regions in particular requires more than descriptions of objects and actions, both of which are what the ML models were trained on. This could likewise influence the fairer and more balanced class-level performance of the Human Labels model, as evidenced by the range of the F1-Scores across all classes. This may also account for why the best-performing models, both overall and class-level, are those that involve human labels.
While Zhang et al. (2021) claim that humans are not necessarily better than machines in choosing discriminative features for image segmentation, our classification model results highlight a different perspective. It is evident that the usage of human labels, whether on its own or in combination with another annotation set, leads to more accurate region classifications. From this, it can be deduced that humans are, in fact, better than machines at creating image annotations that contain these discriminative features in the context of region classification.
Using the Welch’s t-test, we found that there are statistically significant differences between the overall F1-scores from 10 independent runs across the Human Labels model and the ML Objects model at $p$-value $<$ 0.001.
Table 3 shows the Root Mean Square Error (RMSE) for the AdaBoost Regression algorithm used to predict income. Similar to the classification model, the experiment was conducted over 10 independent runs across different combinations of the three annotation sets. The mean and standard deviation of the RMSE values were obtained for the overall model and for each region.
It is apparent that, unlike the classification model, the regression model with ML Objects and ML Captions performed the best overall, indicated by its lowest RMSE value (1279.39). This model also has the smallest range of RMSE values across all classes, following the trend from the best-performing model in the previous experiment, ranging from 168.79 to 2111.71. On the class level, the best-performing models were the ML Objects, ML Object-Captions, and the ML Objects-Human models. However, like the classification model, it is revealed that the models responsible for the best overall and class-level performance are ones that contain ML Objects. Moreover, it is worth highlighting that Asia performs best with the ML Objects model, suggesting that the detected objects in the images from Asia are indicative of income.
The higher performance exhibited by the ML Objects and ML Captions Model is particularly unexpected, especially when contrasted with the trend found in the classification model. However, this observation aligns with the claim made by Zhang et al. (2021) on machines being better at selecting distinguishing features in the context of image segmentation. In our case of income regression, this also applies, although only for ML Objects rather than ML Captions. This could be attributed to certain objects in the images being associated with certain income levels, therefore improving the model’s ability to predict income. Further experiments would need to be conducted to explore this.
Furthermore, we compared the performance differences between the best and worst-performing models (overall), namely the ML Objects-ML Captions Model and the Human Model. The results of the Kolmogorov-Smirnov test showed that the difference is statistically significant ($p$ $<$ 0.001), suggesting that the ML Objects-ML Captions Model’s better performance is statistically reliable.
Figure 4 illustrates a scatter plot of the regression model trained using the ML Objects and ML Captions to predict income on the test dataset. The data points are colour-coded by continent. The blue dotted line represents the line of best fit and the red solid line represents the identity line.
The plot reveals that most of the samples whose incomes are underestimated are mostly countries from Asia, with some samples from the Americas and Europe. On the other hand, most of the overestimated samples are from the Americas, with several samples from Asian and Europe. It is also important to note that none of the African samples deviate significantly from the identity line.
The underestimation of samples from Asia may be attributed to the high number of samples within the lower to middle income range in the training set, resulting in the model skewing predictions for samples from Asia. Interestingly, when looking at the distribution of samples from the Americas, the income levels range from lower to middle income, as well, despite the overestimation by the model. This suggests that perhaps the features detected in the ML Objects and ML Captions for the Americas are not well-represented, causing inaccurate predictions. This is made worse due to the limited number of samples in the dataset, causing there to be less variety in income distribution for particular regions.
Figure 4: Income Regression Model using ML Objects and ML Captions as Annotations Grouped by Continent
Limitations. First, the experiments were done on only one dataset due to the limited availability of image datasets that contain features relating to socio-economic background, such as geographical region and income. Performing the experiments on another similar image dataset would make the results more reliable. Second, the number of image samples in the dataset that have been annotated by humans is rather small. Due to time constraints and limited resources, doing further data collection to obtain more human annotations was not feasible given the scope of this investigation. Having more annotated images would result in better generalisability of the models. Third, due to the small size of the annotated data, under-sampling was not performed for the regression task - unlike the classification model which had under-sampling embedded in the algorithm - which meant that there may be data imbalance when considering number of samples in each region. This could potentially be the reason as to why the performance of some regions is notably better than others. Lastly, there is a risk of misclassification and fault intolerance in this work, as the small dataset could affect the generalisability of the model and, therefore, the trends observed.
Conclusion
Answering the RQs. We first look at RQ1: “How similar are human-generated and ML-generated annotations?”. It can be concluded that from a low-level perspective, i.e., what types of word occur, sentence structures, the ML Objects and ML Captions are similar to each other, but are relatively different to the human labels. This is expected due to underlying cognitive processes during the annotation stage. However, when looking at them from a high-level point of view, despite the apparent low similarity, they are alike in that they exhibit consistent similarity or dissimilarity regardless of the geographical region of the images, indicating lack of regional bias. This led to RQ2: “How do different combinations of annotations affect fairness in ML predictive models?” In the case of region classification, it was revealed that the human labels resulted in the best overall-performing model, as well as the most balanced performance across individual classes. It is additionally apparent that each region obtained best classification performance with models that utilised annotation combinations that included human labels. Similarly, for income regression, the model that has the best overall performance also has the most balanced performance across regions. However, in this case, the combination of ML Objects and ML Captions is what results in this enhanced performance. A similar trend continues in that the best performance for individual classes was obtained with models that contained ML Objects in its annotations. Considering these findings, it can be deduced that there is no particular combination of annotations that result in best overall and most balanced performance as that is highly dependent on the quality of the images and how well the human annotator or machine is able to select the best discriminative features. Additionally, due to imbalanced distributions of income across data samples and low number of samples for certain regions, the regression model is prone to underestimation or overestimation of income for certain regions. Moreover, our analysis suggests that it may be plausible that models with the best overall performance tend to also be more balanced on the class-level, and therefore, fairer. However, further studies need to be done to confirm this.
Implications. These findings have several implications. First, despite their apparent differences, humans and machines may be more alike in their bias, or lack thereof, in the context of visual content. Their consistent similarity and dissimilarity of annotations across regions implies that their level of fairness is comparable. This insight could steer researchers towards further advancing ML models to avoid as much bias as possible. Second, achieving a more balanced performance on the region level can help reduce biases in ML models. This would result in models that not only give accurate outcomes, but ones that are also fair. This would be highly beneficial in real-world scenarios where the model’s outcomes could disproportionately affect certain demographic groups. Third, due to machine annotations (ML Objects and ML Captions) performing better and fairer at the regression task, tasks in other domains could potentially benefit from using machine annotations as well. However, this could result in a potential negative societal impact in that humans become overly reliant on machines, which may cause them to miss possible biases in the models that originated from the training data, miss errors due to assuming the superior performance of the ML models, or even cause difficulties in identifying accountability if the ML models’ prediction errors result in significant repercussions. This could also lead to potential misuse of the technology as it is possible that one intentionally uses a model that is potentially biased without attempting to address the fairness problems.
Acknowledgements
ChatGPT and Claude AI were utilised to assist with brainstorming and to generate sections of this Work, including alternative phrasings and parts of code used for the experiments.